OpenStudy (anonymous):

what percent 0f 60 is 12?

OpenStudy (dmezzullo):

5

OpenStudy (anonymous):

i got 20% is that right?

OpenStudy (anonymous):

So, we have: \[ \frac{12}{60}=\frac{1}{5}=.20=20\% \]So, yes.

OpenStudy (anonymous):

it is 20 percent

OpenStudy (anonymous):

i did it this way.. 60/100 =12 12X 100 = 1200/60 = 20 p = 20%

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

How'd you get \(\frac{60}{100}=12\) ?

OpenStudy (anonymous):

wouldnt it be 12x100/60 then simplify?

OpenStudy (anonymous):

12X100/60 = 1200/60 =20 ?

OpenStudy (anonymous):

yes, and 20 simplified is 5 or 4 because 4x5=20

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

yes got it

OpenStudy (anonymous):

ok, cool :)